# Model-Free Assessment of Simulator Fidelity via Quantile Curves

### v1

@@ 3.2 Calibrated Quantile Curve Theory
which matches the coverage guarantee in Theorem 3.1. We then summarize overall average sim-to-real discrepancy via the calibrated AUC
@@
In words, $\mathrm{AUC}_{\mathrm{cal}}$ aggregates the entire calibrated curve into a single average-bias summary, while $\mathrm{CVaR}^{\mathrm{cal}}_{\alpha}$ averages over the worst $\alpha$ fraction of scenarios under the same finite-sample–corrected quantile indexing.
3.3 Pairwise Simulator Comparison
In Section 3.2, we have developed a distributional characterization of a single simulator’s sim-to-real discrepancy performance. A natural next step is to ask: Can we extend this viewpoint to compare two simulators, i.e., obtain a distributional characterization for pairwise performance?
To that end, we mirror the construction in Section 3.2. Consider two simulators, denoted $S_{1},S_{2}$ and their corresponding point estimates as $\hat{q}^{(1)},\hat{q}^{(2)}$. For a given scenario $\psi$, define the performance discrepancy as
$\delta_{\psi}:=L(p_{\psi},\hat{q}_{\psi}^{(1)})-L(p_{\psi},\hat{q}_{\psi}^{(2)}).$
We are interested in making the inferential claim: $S_{1}$ is at least as good as $S_{2}$, where $\delta_{\psi}<0$ means $S_{1}$ is closer to the ground truth than $S_{2}$ on scenario $\psi$.
The challenge, as in Section 3.2, is that $p_{\psi}$ is unobserved and, with heterogeneous $n_{j}$, the direct plug-in discrepancies $L(\hat{p}_{\psi},\hat{q}_{\psi}^{(1)})-L(\hat{p}_{\psi},\hat{q}_{\psi}^{(2)})$ are non-i.i.d., making it harder to establish uniform claims across quantile levels. We therefore follow the same idea and work with a conservative, data-driven surrogate that constructs a confidence set, enabling a valid distributional comparison between $S_{1}$ and $S_{2}$. The methodology is as follows:
  1. 1.
    For each $j$, compute a confidence set $\mathcal{C}_{j}(\hat{p}_{j})$ on $p_{j}$, and calculate the pseudo-performance discrepancy $\hat{\delta}_{j}:=\sup_{u\in\mathcal{C}_{j}\!\left(\hat{p}_{j}\right)}\{L(u,\hat{q}_{j}^{(1)})-L(u,\hat{q}_{j}^{(2)})\}$.
  2. 2.
    Denote $\hat{U}_{m}(\alpha):=\text{The }\alpha-\text{quantile of }\{\hat{\delta}_{j}\}_{j=1}^{m}$.
With these, we present the following theorem:
Theorem 3.2.
Suppose Assumption 3.1 and 3.2 hold. For any simulation sample size $k\in\mathbb{N}$, denote $\hat{q}^{(1)}_{\psi},\hat{q}^{(2)}_{\psi}$ as point estimates of the two simulators for scenario $\psi$. Define the per-scenario performance discrepancy $\delta_{j}$ and the pseudo-performance discrepancy $\hat{\delta}_{j}$ by
$\delta_{j}$ $:=L\big(p_{j},\hat{q}^{(1)}_{j}\big)-L\big(p_{j},\hat{q}^{(2)}_{j}\big)$
$\hat{\delta}_{j}$ $:=\sup_{u\in C_{j}(\hat{p}_{j},\gamma)}\Big[L\big(u,\hat{q}^{(1)}_{j}\big)-L\big(u,\hat{q}^{(2)}_{j}\big)\Big$],
where confidence set $\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma\right)$ satisfies $\mathbb{P}\big(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma\right)\mid\psi_{j},n_{j}\big)\geq\gamma\quad\text{for a fixed }\gamma\in(0,1).$
Then, for any $\alpha\in(0,1)$ and $\eta\in(0,1)$ such that with probability at least $1-\eta$ over the calibration data $\mathcal{D}$,
$\mathbb{P}_{\psi\sim\Psi}\Big(\delta_{\psi}\leq\widehat{U}_{m}(1-\tfrac{\alpha}{2})\Big|\mathcal{D}\Big)\geq 1-\alpha-\frac{\varepsilon(\alpha,m,\eta)}{\sqrt{m}},$
where
$\varepsilon(\alpha,m,\eta):=$ $\sqrt{2\alpha\log\frac{2m}{\eta}+\frac{\big(\log\frac{2m}{\eta}\big)^{2}+4\log\frac{2m}{\eta}}{m}}$
$+\frac{\log\tfrac{2m}{\eta}+2}{\sqrt{m}}+\sqrt{\frac{\log(4/\eta)}{2}}.$
In particular, the remainder is $O\big(\sqrt{(\log m)/m}\big)$, which vanishes as $m\to\infty$ for any $\alpha$.
The proof of Theorem 3.2 exactly mirrors the argument for Theorem 3.1, where we construct analogous pseudo-discrepancies via confidence sets, and then apply the same quantile-bounding machinery. Intuitively speaking, Theorem 3.2 says that: If the data-driven performance discrepancy quantile $\widehat{U}_{m}(1-\tfrac{\bar{\alpha}}{2})\leq 0$ for some $\bar{\alpha}$, then – up to an $O\big(\sqrt{(\log m)/m}\big)$ adjustment term – $S_{1}$ is at least as good as $S_{2}$ on at least a $(1-\bar{\alpha})$ fraction of scenarios with high probability.
@@ 4.1 Dataset and Methodology
We evaluate our procedure on the WorldValueBench dataset, curated by Zhao et al. (2024) and constructed from the World Values Survey (Haerpfer et al., 2020). The data comprise survey questions asked across 64 countries that elicit attitudes spanning 12 categories (e.g., social values, security, migration). Individual-level covariates are available for each respondent, including gender, age, migration status, education, marital status, etc.
Figure 2: Example of World Value Questions. Retrieved from Haerpfer et al. (2020).
@@
After data cleaning, we retain 235 distinct questions and responses from 96,220 individuals. Each question offers a categorical set of possible answers, as illustrated in Figure 2. To place heterogeneous answer sets on a common scale, we map each question’s categories to the interval [1,1][-1,1], producing a bounded real-valued outcome for every scenario. Individual-level covariates are then used to construct synthetic profiles and corresponding prompts for simulation. Additional details on preprocessing and dataset characteristics are provided in Appendix C.
@@
We generate synthetic responses and estimate $\{(\hat{q}_{j})\}_{j=1}^{235}$ for four LLMs: GPT-4o (gpt-4o), GPT-5 mini (gpt-5-mini), Llama 3.3 70B (Llama-3.3-70B-Instruct-Turbo), and Qwen 3 235B (Qwen3 235B A22B Thinking 2507 FP8). As a benchmark, we also construct a Uniform baseline: for each question, this generator samples an answer uniformly at random from the available choices for each individual. We set the simulation budget to $k=500$ for each question and then randomly select 200 of these simulated responses to estimate $\hat{q}_{j}$. The same 500 human participants used to generate the synthetic profiles are also used to calculate $\{\hat{p}_{j}\}_{j=1}^{235}$ and construct the corresponding confidence sets $\mathcal{C}_{j}$. Because some participants decline or are unable to answer certain questions, the effective human sample size $n_{j}$ to construct $\hat{p}_{j}$ varies slightly across questions, typically ranging from 450 to 500.
Finally, since outcomes are bounded within [1,1][-1,1], we construct per-question confidence sets $\mathcal{C}_{j}$ following Example 3.2, and compute the corresponding pseudo-gaps under our general framework. We then apply the procedure described in Section 3 to obtain a fidelity profile for each candidate LLM. Throughout, we use the squared-error discrepancy $L(p,q)=(p-q)^{2}$, and set the confidence-level parameter $\delta=0.05$. For completeness, we also apply the same workflow to a Bernoulli setting (EEDI) and a multinomial setting (OpinionQA); see Appendix E.
@@ 4.2 Fidelity Profiling
Figure 3: Calibrated $V(\alpha)$ across LLMs.
[…]
Figure 4: Robustness check of simulator performance under different $n$-levels.
We also provide a robustness check on whether this dominance of performance is an anomaly for this particular set of numbers of $n_{j}$. Figure 4 shows performance across LLMs given a fixed $n_{j}$, where we consider $n_{j}=\{50,500,5000,10000\}$. The results indicate that the dominance of GPT-4o over GPT-5-mini, Llama 3.3 70B, and QWen-3-235B is consistent across $n_{j}$, hence validating our method’s result.
@@ 5 Tightness Analysis of Calibrated Quantiles
A natural concern with our approach is the tightness of the calibrated quantile curve from Theorem 3.1. Since we rely on off-the-shelf concentration inequalities to construct pseudo-gaps and empirical distribution gaps, it might be the case that our method returns a very loose quantile that is of little practical use. In order to address this concern, we conduct numerical analysis to isolate the effect of such finite-sample calibration and order-statistic adjustments from the simulation misalignment. We run the following analysis:
Since we have 96,22096{,}220 respondents per question in the World Value Bench dataset, we treat the estimate $\hat{p}_{j}$ based on all samples as the underlying true $p_{j}$. Therefore, we can calculate the oracle gap $\Delta_{j}^{*}:=L(p_{j},\hat{q}_{j})$ for each question $j$, and hence construct the true quantile function. We set $k=200$ and randomly sample $n_{j}=100,200,500,1000$ to calculate $\{\hat{\Delta}_{j}\}$ under different finite-sample calibration and order-statistic adjustments to evaluate the tightness of our proposed calibrated quantiles. We present our result in Figure 5.
@@
Figure 5: Tightness analysis of different $n_{j}$ under GPT-4o.
As Figure 5 shows, the envelope quantile is loose when we only have $n_{j}=100$. Nonetheless, it converges to the true quantile $\{\Delta^{*}\}$ when $n$ grows to 200, 500, 1,000, indicating that our method is tight given sufficient real-world sample size per scenario.
However, in most cases we will not have such a large dataset as benchmark to test tightness. Subsequently, we will provide a data-driven tightness guarantee.
Theorem 5.1.
@@
Suppose the setup of Theorem 3.1 and Assumptions 3.13.2 hold. Let $\gamma\in(1/2,1]$ and define lower pseudo-discrepancies
$\Delta_{j}^{-}:=\inf_{u\in C_{j}(\hat{p}_{j})}L(u,\hat{q}_{j}),$
where $C_{j}(\hat{p}_{j})\subset\Theta$ are data-driven confidence sets satisfying $\mathbb{P}\big(p_{j}\in C_{j}(\hat{p}_{j})\mid\psi_{j},n_{j}\big)\geq\gamma.$ Let $\hat{V}_{m}^{-}(\alpha)$ denote the empirical $\alpha$-quantile of $\{\Delta_{j}^{-}\}_{j=1}^{m}$, and $\hat{V}_{m}(\alpha)$ the empirical $\alpha$-quantile of the upper pseudo-discrepancies from Theorem 3.1 (under the same coverage level $\gamma$).
For any $\eta\in(0,1)$, the following uniform lower bound holds: with probability at least $1-\eta$ over the calibration data $\mathcal{D}$, for all $\alpha\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}\big(\gamma\alpha\big)\Big|\mathcal{D}\Big)\geq 1-\alpha-\frac{\varepsilon_{-}^{(\gamma)}(\alpha,m,\eta)}{\sqrt{m}},$ (5.1)
where $\varepsilon_{-}^{(\gamma)}(\alpha,m,\eta)$ is of order $O\big(\sqrt{\tfrac{\log m}{m}}\big)$ uniformly in $\alpha$.
Let $V:[0,1]\to\mathbb{R}$ denote the quantile function of $\Delta(\psi)$. Then, for any fixed $\tau\in(0,1)$, as $m\to\infty$ we obtain the asymptotic band
$\hat{V}_{m}^{-}\big(\gamma\tau\big)\leq V(\tau)\leq\hat{V}_{m}\big(\gamma+(1-\gamma)\tau\big)+o_{m}(1),$ (5.2)
where $o_{m}(1)\to 0$ as $m\to\infty$.
Proof is deferred to Appendix B. The main takeaway is that by adopting Theorem 5.1 and  3.1, we can calibrate a confidence band for the actual quantile curve.
Remark 2.
As discussed in Section 3.1, the factor $\gamma$ in (5.1) and (5.2) directly reflects the per-scenario coverage level of the confidence sets $C_{j}(\hat{p}_{j})$. For the tail behavior (e.g., small $\tau$), the upper side of the band is intrinsically conservative: we only guarantee
$V(\tau)\leq\hat{V}_{m}\big(\gamma+(1-\gamma)\tau\big)+o_{m}(1),$
so when $\tau$ is close to 0 the index $\gamma+(1-\gamma)\tau$ is bounded away from $1$. In other words, the right-hand side controls $V(\tau)$ using a substantially more central empirical quantile, and the resulting upper band can be quite loose in the extreme left tail. The same phenomenon appears symmetrically in the right tail when $\tau\approx 1$. For any fixed $\gamma<1$, this conservativeness near very small or very large $\tau$ is unavoidable: the band must widen in the extremes to pay for the finite per-scenario coverage level.
@@ 6 Discussion
We present a model-free estimator for the quantile function that makes no parametric assumptions on either the simulator or the ground truth, delivers finite-sample guarantees for any desired $\alpha$-quantile, and supports flexible summaries from means to tail-risk measures. We further develop a pairwise comparison framework to directly compare simulators’ ability to replicate real-world outcomes. Finally, we apply our methods to a real-world dataset, assessing multiple LLMs and demonstrating their relative effectiveness at simulating human responses.
Despite its broad applicability, the method we introduce leaves several natural extensions. First, our proof relies on DKW-type concentration – often conservative for small $m$ – and a grid-uniform step that further loosens constants; tightening these bounds is an immediate target. Second, many applications involve temporally dependent, dynamic simulation processes; extending our static framework to dynamic settings would broaden applicability. Third, our analysis assumes i.i.d. scenarios, whereas covariate shift or endogenous sampling may invalidate marginal guarantees; addressing such distribution shifts is an important avenue for future work.

### v2

@@ 3.3 Benefits of Distribution-Level Characterization
which matches the coverage guarantee in Equation 3.4 in asymptotic regime. We then summarize overall average sim-to-real discrepancy via the calibrated AUC
@@
In words, $\mathrm{AUC}_{\mathrm{cal}}$ aggregates the entire calibrated curve into a single average-bias summary, while $\mathrm{CVaR}^{\mathrm{cal}}_{\alpha}$ averages over the worst $\alpha$ fraction of scenarios under the same index-adjusted quantile curve.
@@ 4.1 Dataset and Methodology
We evaluate our procedure on the WorldValueBench dataset, curated by Zhao et al. (2024) and constructed from the World Values Survey (Haerpfer et al., 2020). The dataset comprises survey questions asked across 64 countries, covering 12 thematic categories (e.g., social values, security, migration). Individual-level covariates are available for each respondent, including gender, age, migration status, education, marital status, and related demographics.
Figure 2: Example of World Value Question. Source: Haerpfer et al. (2020).
@@
After data cleaning, we retain 235 distinct questions and responses from 96,220 individuals. Each question offers a categorical set of possible answers, as illustrated in Figure 2. To place heterogeneous answer sets on a common scale, we map each question’s categories to the interval [1,1][-1,1], producing a bounded real-valued outcome for every scenario and respecting the natural ordering of response options when available (e.g., disagree–neutral–agree). Individual-level covariates are then used to construct synthetic profiles and corresponding prompts for simulators. Additional details on preprocessing and dataset characteristics are provided in Appendix C.
@@
We generate synthetic responses and estimate $\{\hat{q}_{j}\}_{j=1}^{235}$ for four LLMs: GPT-4o (gpt-4o), GPT-5 mini (gpt-5-mini), Llama 3.3 70B (Llama-3.3-70B-Instruct-Turbo), and Qwen 3 235B (Qwen3 235B A22B Thinking 2507 FP8). As a benchmark, we also construct a uniform baseline: for each question, this generator samples an answer uniformly at random from the available choices for each individual. We simulate $200$ LLM samples per question, and use these simulated responses to estimate $\hat{q}_{j}$. For humans, we subsample 500 respondents per question to calculate $\{\hat{p}_{j}\}_{j=1}^{235}$ and construct the corresponding confidence sets $\mathcal{C}_{j}$. Due to nonresponses, the effective sample size $n_{j}$ varies across questions (typically 450–500).
Since the outcome is bounded in [1,1][-1,1], we construct per-scenario confidence sets $\mathcal{C}_{j}$ using Example 3.2. We set $\gamma_{j}=1-n_{j}^{-1/3},$ which is increasing in sample size $n_{j}$ [^2] As discussed in Section 3, the coverage level $\gamma_{j}=g(n_{j})$ should increase with the sample size $n_{j}$. We therefore adopt a generic form $\gamma_{j}=1-n_{j}^{-\beta},$ and conduct a sensitivity analysis on $\beta$ value in Appendix C.4., and yields an average coverage level of $\bar{\gamma}_{m}\approx 0.87$. The loss function $L(p,q)=(p-q)^{2}$. Next, we use the procedure described in Section 3 (see (3.1)-(3.3)) to obtain a fidelity profile for each candidate LLM, and plot the calibrated curve using the asymptotic index adjustment in (3.5). For completeness, we also apply the same workflow to a Bernoulli setting (EEDI) and a multinomial setting (OpinionQA); see Appendix E.
@@ 4.2 Fidelity Profiling
Figure 3: Calibrated $\hat{V}(\tau)$ across LLMs.
The calibrated quantile functions for the four different LLMs are plotted in Figure 3. Lower, flatter curves indicate uniformly small discrepancies, while elbows reveal rare but severe misses. Our results suggest that all LLM-simulators outperform the uniform benchmark on over $80\%$ of questions, but fail to capture the tail scenarios. GPT-4o is the lowest across all quantiles, indicating the most reliable alignment, with GPT-5-mini close behind yet failing to capture some outlier questions, and Llama 3.3 and QWen-3 trail behind.
Figure 4: Robustness check of simulator performances.
We also provide a robustness check on whether the dominance relation is an anomaly of the particular choice for the sample size $n_{j}$. Figure 4 shows performance of the LLMs for $n_{j}=\{50,500,5000,10000\}$. The results indicate that the dominance relation between LLMs remains the same.
@@ 5 Tightness Analysis of Calibrated Quantiles
A natural concern with our approach is the tightness of our calibrated quantile curve. Since our construction relies on off-the-shelf concentration bounds and worst-case pseudo-discrepancies, the resulting calibrated quantiles could, in principle, be conservative to the point of being vacuous. To assess this empirically, we conduct a tightness study in a setting that leverages a feature of WorldValueBench.
We leverage the WorldValueBench benchmark, which contains up to 96,22096{,}220 respondents per question. We treat the empirical distribution $\hat{p}_{j}$ computed from all available respondents as a proxy for the true human response distribution $p_{j}$. This enables computation of an “oracle” discrepancy $\Delta_{j}^{\star}:=L(p_{j},\hat{q}_{j}),$ and hence an oracle quantile curve across questions. We fix the simulator budget at $k=200$ and, for each question $j$, subsample $n_{j}\in\{50,500,1000,5000\}$ respondents to construct the confidence sets and pseudo-discrepancies $\{\hat{\Delta}_{j}\}_{j=1}^{m}$. We then compare the resulting calibrated curves against the oracle quantile curve.
Figure 5: Tightness analysis for varying $n_{j}$ under GPT-4o.
The results plotted in Figure 5 indicate that the calibrated envelope is substantially conservative when $n_{j}=50$, but becomes tighter when $n_{j}\geq 500$. This behavior is consistent with our construction: larger $n_{j}$ yields more concentrated estimates $\hat{p}_{j}$ and smaller confidence sets $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$, which reduces the inflation induced by the supremum defining $\hat{\Delta}_{j}$.
The results in Figure 5 correspond to $\gamma_{j}=1-n_{j}^{-\frac{1}{3}}$. In order to isolate the impact of batch-size-specific $\gamma_{j}$, as we discussed in Section 3, we repeat the same experiment by setting $\gamma_{j}\equiv\tfrac{1}{2}$ for all questions. The corresponding results are plotted in Figure 6. The plots here exhibit a persistent vertical separation between the oracle quantile curve and the calibrated curves, even for large $n_{j}$. This phenomenon is explained by the fact that here $\bar{\gamma}_{m}=\tfrac{1}{2}$ (independent of $n_{j}$) and the asymptotic calibration evaluates the pseudo-quantile at $\hat{V}_{m}\big(0.5\tau+0.5\big)$. Consequently, at lower quantile levels ($\tau\approx 0$), the our procedure approximates the quantile function by at least the median, inducing a non-vanishing conservativeness that cannot be eliminated by increasing $n_{j}$. This underscores the importance of allowing scenario-dependent coverage levels.
@@
Figure 6: Tightness analysis for varying $n_{j}$ under GPT-4o with fixed $\gamma_{j}=\tfrac{1}{2}$.
While WorldValueBench provides a large pool of respondents that facilitates an oracle tightness check, most applications do not enjoy such extensive ground-truth data. We next develop a complementary tightness assessment: a confidence band that sandwiches the oracle quantile curve.
Theorem 5.1 (Informal Version).
@@
Suppose the setup of Theorem 3.1 and Assumptions 3.13.2 hold. Let $\{\gamma_{L,j},\gamma_{U,j}\}_{j=1}^{m}\subset(0,1]$ be coverage levels and define the lower and upper pseudo-discrepancies
$\hat{\Delta}_{j}^{-}:=\inf_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{L,j})}L(u,\hat{q}_{j}),\ \hat{\Delta}_{j}^{+}:=\sup_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{U,j})}L(u,\hat{q}_{j}),$
where $\mathcal{C}_{j}(\hat{p}_{j},\gamma)\subset\Theta$ are $\gamma$-confidence sets. Let $\hat{V}_{m}^{-}(\cdot)$ (resp. $\hat{V}^{+}(\alpha)$) denote the empirical quantile function of $\{\hat{\Delta}_{j}^{-}\}_{j=1}^{m}$ ( resp. $\{\hat{\Delta}_{j}^{+}\}_{j=1}^{m}$). Let $V:[0,1]\to\mathbb{R}$ denote the quantile function of $\Delta(\psi)$. Then, as $m\to\infty$, for any fixed $\tau\in(0,1)$,
$\hat{V}_{m}^{-}\big(\bar{\gamma}_{L,m}\tau\big)\lesssim V(\tau)\lesssim\hat{V}_{m}^{+}\big(\bar{\gamma}_{U,m}\tau+(1-\bar{\gamma}_{U,m})\big),$ (5.1)
where $\bar{\gamma}_{L,m}=\tfrac{1}{m}\sum_{j=1}^{m}\gamma_{L,j}$, and $\bar{\gamma}_{U,m}=\tfrac{1}{m}\sum_{j=1}^{m}\gamma_{U,j}$.
Figure 7: Confidence Bands of GPT-4o and Llama.
The formal version of this result is presented in Theorem B.1 and its proof is deferred to Appendix B.4. The takeaway is that Theorems 3.1 and B.1 provide complementary tools for assessing simulator fidelity at the distributional level. Theorem 3.1 provides a one-sided, calibrated upper envelope for the population quantile curve, while Theorem B.1 augments this with a lower envelope, producing a confidence band that sandwiches the unknown quantile function $V(\tau)$. Figure 7 illustrates these bands for two subsample sizes $n_{j}\in\{1000,5000\}$ and compares GPT-4o against Llama 3.3. Across both panels, the GPT-4o band lies systematically below the Llama 3.3 band over most quantile levels, with the separation becoming most pronounced in the upper tail. This pattern indicates that GPT-4o achieves systematically smaller calibrated discrepancy across scenarios and exhibits fewer extreme failures, and the conclusion is stable as $n_{j}$ increases. Importantly, the comparison is fully data-driven: unlike oracle evaluations that rely on near-population human samples, the band construction does not require access to ground-truth discrepancies and can be carried out using only the observed heterogeneous batches.
@@ 6 Discussion
We develop a model-agnostic framework for profiling simulator fidelity through the quantile function of the sim-to-real discrepancy. Our approach makes no parametric assumptions about either the simulator or the real system, rigorously accounts for finite and heterogeneous sampling across scenarios, and yields calibrated, finite-sample guarantees at any target quantile level. The resulting fidelity profile supports a range of downstream summaries, from average performance to tail-risk measures that isolate rare but consequential failures. Finally, we demonstrate the practical value of the framework on real survey data, evaluating multiple LLM-based simulators and illustrating systematic, distribution-level differences in their alignment with human populations.
Despite its broad applicability, the method we introduce leaves several natural extensions. First, our proof relies on DKW-type concentration – often conservative for small $m$ – and a grid-uniform step that further loosens constants; tightening these bounds is an immediate priority. Second, many applications involve temporally dependent, dynamic simulation processes; extending our static framework to dynamic settings would broaden applicability. Third, our analysis assumes i.i.d. scenarios, whereas covariate shift or endogenous sampling may invalidate marginal guarantees; addressing such distribution shifts is an important avenue for future work.